A rotating beacon is located a distance d from a long wall. The distance d is given by
d = 3 tan 2πt,
where t is the time measured in seconds since the beacon started rotating. Solve the equation for t.

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

d = 3 tan 2πt,
where t is the time measured in seconds since the beacon started rotating. Solve the equation for t.
-------
tan(2pit) = d/3
------
2*pi*t = invtan(d/3)
-------
t = (1/(2pi))*invtan(d/3)
------------
Cheers,
Stan H.
----------------

Answer by ikleyn(52855)

.
It seems to me that the problem in whole doesn't make sense.

There is no connection between the formula and rotating beacon. (and rotating at all).